OpenStudy (anonymous):

The length of your rectangular backyard is 275 feet and the width is 125 feet. You want to enclose the backyard with chain link fence that is $1.25 per foot. Calculate the total cost of the fencing required to enclose the backyard.

OpenStudy (anonymous):

275•125=?

OpenStudy (anonymous):

With that product, you multiply it by $1.25

OpenStudy (anonymous):

uhhh i got 35,234,374

OpenStudy (anonymous):

ion think thats right

OpenStudy (anonymous):

275•125=34,375 square feet 34,375•1.25=$41,250

OpenStudy (aaronq):

They're asking for the perimeter not the area. Perimeter = 2L + 2W = 2(275 feet) + 2(125 feet) Cost = Perimeter * price = [2(275 feet) + 2(125 feet)]*($1.25/foot) = $ 1000